### v1

# Speculative End-Turn Detector for Efficient Speech Chatbot Assistant

Abstract
Spoken dialogue systems powered by large language models have demonstrated remarkable abilities in understanding human speech and generating appropriate spoken responses. However, these systems struggle with end-turn detection (ETD)—the ability to distinguish between user turn completion and hesitation. This limitation often leads to premature or delayed responses, disrupting the flow of spoken conversations. In this paper, we introduce the ETD Dataset, the first public dataset for end-turn detection. The ETD dataset consists of both synthetic speech data generated with text-to-speech models and real-world speech data collected from web sources. We also propose SpeculativeETD, a novel collaborative inference framework that balances efficiency and accuracy to improve real-time ETD in resource-constrained environments. Our approach jointly employs a lightweight GRU-based model, which rapidly detects the non-speaking units in real-time on local devices, and a high-performance Wav2vec-based model running on the server to make a more challenging classification of distinguishing turn ends from mere pauses. Experiments demonstrate that the proposed SpeculativeETD significantly improves ETD accuracy while keeping the required computations low. Datasets and code will be available after the review.
1 Introduction
Recent advancements in large language models (LLMs) have spurred extensive research into various LLM-based agents (Song et al., 2023; Shao et al., 2024; Zhang et al., 2024; Xie et al., 2024). Many studies focus on integrating LLMs with spoken dialogue systems to enable more engaging human-computer interaction (Mitsui et al., 2023; Yan et al., 2024; Ma et al., 2024; Veluri et al., 2024). However, LLM-based spoken dialogue systems face a critical challenge: when the user suddenly stops speaking, the systems often fail to discern whether the user has completed their turn or is merely pausing to think during a conversation (Lin et al., 2022; Umair et al., 2024). For instance, as illustrated in Figure 1, when a user momentarily pauses while contemplating a specific question, the LLM may erroneously interpret this pause as the end of the turn, leading to premature and off-target responses. This task, called end-turn detection (ETD), is a challenging task that requires nuanced mechanisms based on a firm understanding of the content of the speech; naïve cues, such as the pause duration, fails to provide meaningful information to distinguish between the turn completion and user hesitation (Ten Bosch et al., 2005).
Recent research has explored techniques to improve turn-taking in dialogue systems to address this issue (Chang et al., 2022; Veluri et al., 2024). However, a major bottleneck is the lack of publicly available datasets. Existing datasets, such as the Fisher corpus (Cieri et al., 2004), are now paid for utilizing, while studies like Chang et al. (2022); Veluri et al. (2024); Ma et al. (2024) rely on private, inaccessible data.
To tackle this bottleneck, we construct and release the ETD dataset, an open-to-public dataset that is specifically designed for training and evaluating end-turn detection models. The ETD dataset consists of more than 120k samples with over 300 hours of conversational data. In particular, the dataset consists of both synthetic spoken conversation generated via applying text-to-speech (TTS) generation on text dialogue dataset and real-world dialogues collected from web sources.
| Model | Param. | Acc. |
| GRU | M | 79.34 |
| Wav2vec 2.0 | M | 99.49 |
Beyond dataset construction, we propose an efficient end-turn detection algorithm tailored for real-time chatbot interactions under resource-constrained environments. While transformer-based audio processing models achieve strong performance, their high computational cost makes them impractical to be used in highly repetitive tasks with real-time applications, such as ETD. In contrast, models under 1M parameters can run efficiently on local devices (Qi et al., 2023; Wu & Chen, 2023) but suffer from degraded accuracy (see Table 1). To address this gap, we introduce SpeculativeETD, a novel collaborative inference framework that balances efficiency and accuracy. In this framework approach, we adopt both the lightweight GRU-based model (Chung et al., 2014) and the high-performance Wav2vec-based model (Baevski et al., 2020). Here, the lightweight model operates on local devices in real time and keeps track of whether the speaker is speaking or not (either as a pause or end-of-turn). Whenever the model detects the silence, the model queries the high-performance model to make a more challenging prediction, i.e., telling whether the silence marks the end of the speaker’s turn or not. This framework removes the need for the high-performance model to be running in real-time, dramatically saving the amount of required computations with minimal degradation in accuracy.
Our key contributions can be summarized as follows:
We release the first open-source dataset for end-turn detection, incorporating synthetic and real-world speech data.
We propose a novel method, SpeculativeETD, for efficient and accurate end-turn detection, leveraging a lightweight on-device model and a server-side model.
2 Related work
Turn-taking.
Turn-taking plays a crucial role in human communication, allowing speakers to transition smoothly without interruptions. Its detection is guided by linguistic, prosodic, and non-verbal cues such as pitch variation, speech rhythm, and gaze direction (Gravano & Hirschberg, 2011; Levinson & Torreira, 2015). Early turn-taking models have used finite-state machines to predict turn timing and duration (Raux & Eskenazi, 2009).
In spoken dialogue systems, turn-taking ensures fluid human-computer interactions. Various approaches have been explored for turn boundary prediction, including end-to-end models based on automatic speech recognition (ASR) (Chang et al., 2022) and transformer-based models incorporating contextual and pragmatic cues (Ekstedt & Skantze, 2020). More recently, full-duplex LLM-based agents have been developed to enable simultaneous listening and responding (Défossez et al., 2024). However, existing models remain computationally expensive and rely on private datasets, limiting real-time applicability and reproducibility. Our approach addresses these challenges by introducing an efficient turn-detection method and an open-source dataset.
Disfluency Detection.
Disfluencies, such as interruptions, filler words, or self-corrections, occur frequently in spontaneous speech and can disrupt the natural flow of conversations. Identifying and handling these phenomena is crucial for maintaining coherent interactions in spoken dialogue systems. Traditional approaches primarily rely on text-based detection, analyzing transcriptions with span classification models to improve accuracy (Ghosh et al., 2022). Although effective, these methods depend on ASR, which introduces latency and errors. Recent studies have explored direct speech-based disfluency detection (Zhou et al., 2024) and integrated frameworks that combine ASR with disfluency removal for smoother real-time processing (Lou & Johnson, 2020). However, these methods often struggle with spontaneous speech variations and real-time efficiency.
Chatbot Agents.
LLMs have advanced chatbot systems, enabling more natural spoken dialogue interactions (Lin et al., 2022; Rubenstein et al., 2023). Traditional chatbots rely on turn-based exchanges, whereas recent models explore real-time processing to handle overlapping speech, interruptions, and backchanneling (Veluri et al., 2024; Ma et al., 2024).
However, LLM-driven agents often suffer from high computational costs due to their autoregressive inference paradigm, leading to slow response times. Additionally, while some models aim to adaptively handle turn-taking, they frequently miss fine-grained conversational cues like hesitations and mid-turn pauses, resulting in suboptimal interactions (Umair et al., 2024). Another challenge is the reliance on proprietary datasets, restricting accessibility for further improvements (Cieri et al., 2004). To address these challenges, we incorporate speculative inference for efficient real-time processing and introduce an open-source dataset for end-turn detection, fostering further research in spoken dialogue systems.
3 End-Turn Detection (ETD) Dataset
Before introducing the dataset, we first formally describe the task of end-turn detection (ETD). The goal of ETD is to infer from the given speech whether the speaker has finished speaking or not. More precisely, at each time , the speaker is at one of the three states. (1) Speaking unit (SU): The speaker is in speech. (2) Pause: The speaker is not in speech but is only at a brief pause and intends to keep on talking. (3) Gap: The speaker has finished talking, marking the end of his or her turn.
The ETD dataset consists of multiple audio speech files labeled with the ternary segmentation information. That is, the label of each audio file is a sequence of the (state, starting time) pair of each same-state segment, e.g.,
We construct the ETD dataset using both synthetic (for both training and evaluation) and real (for evaluation only) conversational audio segments.
3.1 Synthetic data generation
We have constructed the synthetic audio dataset of multi-turn conversation by applying the TTS (text-to-speech) model to the existing text dialogue dataset. In particular, we have used the MultiWOZ corpus (Budzianowski et al., 2018) as the base dataset. As the TTS model, we have used multiple TTS models, in particular the MeloTTS (Zhao et al., 2023) and Google Cloud Text-to-Speech, to ensure the diversity of the generated speech.
An important thing to note here is that a naïve text-to-speech translation of text dialogues will generate data consisting only of speaking units and gaps, as the textual cues for the “pause” are typically missing in a text dialogue. Thus, we artificially insert the pauses to the conversation in three different styles (Figure 2(a)):
Base: We do not insert any pause.
w/ Pause: We randomly select tiny hesitation—typically inserted by TTS models to mimic human breathe (Braunschweiler & Buchholz, 2011; Hwang et al., 2023; Yang et al., 2024)—and extend it into a pause. The duration of the pause has been randomly drawn from the distribution (where the unit is in seconds). For some randomly picked samples, we have also removed the segments that follow the pause, generating samples without a gap at the end of the audio file.
w/ Filler words: We randomly inject the filler words, e.g., “um” or “uh,” at arbitrary locations in the text dialogue. After the TTS generation, we add a pause after the filler words. Again, we randomly select a fraction of samples and remove the segments after the pause to generate a sample that does not end with a gap.
3.2 Real data generation
We have also collected real spoken conversation data from YouTube and Buckeye speech corpus (Pitt et al., 2005); we have selected these sources as they contain many spontaneous, natural dialogues between two speakers with frequent turn-takings. To ensure that the collected samples involve exactly two speakers, we have applied a pretrained speaker diarization toolkit (Bredin, 2023) to each audio file and filtered out those with more than two identified speakers.
Given conversation data, we again utilized the speaker diarization tool to generate segmentations of the speech of two different speakers (Figure 2 (b)). Any silent interval between speech segments that exceed 200ms is labeled as either pause or gap, depending on whether the neighboring speech segments belong to the same speaker or not; here, the 200ms criterion follows the commonly used thresholds in the turn-taking literature (Stivers et al., 2009; Heldner & Edlund, 2010; Nguyen et al., 2022; Li et al., 2022). Similarly to the synthetic dataset, we have randomly selected some samples and removed the audio segment that follows a randomly chosen gap or pause.
3.3 Other details
Splits and statistics. The default train/dev/test split of the synthetic data follows that of the MultiWOZ dataset (Budzianowski et al., 2018). The statistics of the synthetic and real datasets are given in Table 8 and Table 9, respectively.
Files. For real data crawled from YouTube, we do not directly share the audio files to avoid license issues. Instead, we release the code and public URL to download the audio files.
| All | V1 (base) | V2 (w/ pause) | V3 (w/ filler words) | |
| Total | ||||
| # Samples | 122,481 | 28,368 | 47,076 | 47,037 |
| # Total duration (h) | 200.86 | 40.91 | 89.74 | 70.21 |
| # Average duration (s) | 5.90 | 5.19 | 6.87 | 5.37 |
| Train | ||||
| # Samples | 96,773 | 22,321 | 37,246 | 37,206 |
| # Total duration (h) | 158.4 | 32.14 | 70.8 | 55.46 |
| # Average duration (s) | 5.89 | 5.18 | 6.84 | 5.37 |
| Dev | ||||
| # Samples | 12,840 | 3,008 | 4,916 | 4,916 |
| # Total duration (h) | 21.25 | 4.34 | 9.48 | 7.43 |
| # Average duration (s) | 5.96 | 5.20 | 6.94 | 5.44 |
| Test | ||||
| # Samples | 12,868 | 3,039 | 4,914 | 4,915 |
| # Total duration (h) | 21.21 | 4.43 | 9.46 | 7.32 |
| # Average duration (s) | 5.93 | 5.24 | 6.93 | 5.36 |
| All | {SU,Pause} | {SU,Gap} | {SU,Pause,Gap} | Total Pause | Total Gap | |
| Total | ||||||
| # Samples | 8,022 | 3,283 | 1,369 | 3,370 | - | - |
| Duration (h) | 115.08 | 13.56 | 2.64 | 98.88 | 16.09 | 12.14 |
| YouTube | ||||||
| # Samples | 6,357 | 2,858 | 1,179 | 2,320 | - | - |
| Duration (h) | 82.84 | 10.30 | 2.28 | 70.26 | 9.76 | 8.90 |
| Buckeye | ||||||
| # Samples | 1,665 | 425 | 190 | 1,050 | - | - |
| Duration (h) | 32.25 | 3.26 | 0.36 | 28.68 | 6.33 | 3.24 |
4 Method
In this section, we describe SpeculativeETD, a framework that achieves a favorable tradeoff of computational efficiency and accuracy for end-turn detection.
For a smooth real-time conversation, one needs to run the end-turn detector repetitively at a rapid frequency. Thus, it is desirable to deploy an ETD model with an extremely small computational footprint. However, lightweight models—e.g., a recurrent model with a single GRU layer—tend to suffer from a low detection accuracy. On the other hand, heavier transformer-based models typically require a large computational cost despite being highly accurate in the end-turn detection (see Section 6).
To enjoy the best of both worlds, the SpeculativeETD is designed as a collaborative inference framework that utilizes both a lightweight model (say, GRU), which runs on-device, and a heavier model (say, Wav2vec 2.0), which runs on the server. In a nutshell, the proposed SpeculativeETD operates in two stages (Figure 3):
1.
On-device: The lightweight recurrent model processes the streaming audio signal frame-by-frame (or chunk-by-chunk) and detects whether the user’s speech has finished. In other words, the model conducts a binary classification between the Speaking Unit (SU) and non-SU (Gap and Pause). As this task is relatively easier than distinguishing between gaps and pauses, we can achieve high accuracy with tiny computational costs and latency.
2.
Server-side: Given a speech segment that ended with a silence, the heavier transformer-based model (e.g., a Wav2Vec 2.0) predicts whether the silence indicates the gap or the pause. Whenever the gap is detected, the language model will be called for generating an appropriate response to the speaker.
The name of the method—SpeculativeETD—comes from its structural resemblance to the speculative decoding (Leviathan et al., 2023), a popular decoding algorithm for large language models that use a smaller and faster drafter model for autoregressive generation and larger models for parallel verification. We note, however, that our method critically differs from the speculative decoding in the sense that the small model and large model do not predict the same output classes in our framework; rather than playing the role of a verifier, the large model is used for making a more difficult fine-grained prediction, conditioned on the predictions of the small model.
Advantages.
The key advantages of this framework are twofold: (1) The trickier decision that requires a large computation (i.e., telling “Gap” vs. “Pause”) can happen only once per each consecutive segment of silence, not at every time frame. As we will see in Section 6, this helps save the computation more than 10 with only a small degradation prediction accuracy. (2) The two-stage pipeline allows us to utilize on-device computation by placing the lightweight model on the edge. By doing so, we no longer need the communication between the edge device and the server to be done in a continuous manner; instead, we can communicate once every silence.
Inference Protocol.
During real-time operation, the on-device model processes each new audio frame in constant time. Predictions are queued locally, and when predicted as ”not speaking,” they are sent to the server. The server-side model then processes the segment, determining whether an end-turn event has occurred or if streaming should continue.
5 Experimental Setup
We now empirically validate the usefulness of the constructed ETD dataset (Section 3) and the effectiveness of the proposed method, SpeculativeETD (Section 4).
Task. We consider two scenarios of the ETD task:
Binary classification: When a speech segment comes as an input, the detector reads the whole audio and determines whether the speech segment has ended (Gap) or not (Pause).
Real-time audio segmentation: The speech arrives in real-time, and at each interval (e.g., every 100ms), the detector determines whether the current state of the speaker is SU, Pause, or Gap. That is, we are conducting a ternary segmentation in a sequential manner.
Models.
We evaluate a total of four models for end-turn detection.
VAP (Ekstedt & Skantze, 2022): A popular open-source pretrained turn-taking model, which can also be employed for end-turn detection tasks. We use the pretrained encoder as frozen and train the predictor module of the model from scratch using the ETD dataset. We have used the “comparative” predictor head, which fits the ETD task.
GRU (Chung et al., 2014): We construct a lightweight GRU model with 1M parameters, with two Conv2D layers and one GRU layer, and train from scratch with the ETD dataset.
Wav2vec 2.0 (Baevski et al., 2020): A relatively large transformer-based model with 94M parameters, which achieves high performance in many speech processing tasks. We fine-tune the full model using the ETD dataset.
SpeculativeETD (Ours): We construct an ETD model using the GRU and Wav2vec 2.0 model following the same settings as described above. Note that, as this method assumes a real-time scenario, we evaluate this method only in real-time audio segmentation scenarios and not in binary classification.
Evaluation.
For the binary classification task, performance is quantified using Precision, Recall, F1-score, and accuracy metrics. In the real-time segmentation task, we assess the F1-score at 100ms intervals and evaluate segmentation quality using the Intersection over Union (IoU) metric for the average of each of the three classes (Speaking Unit, Gap, Pause).
Datasets for training/evaluation.
In all experiments, we have trained the models using the training split of the synthetic data in the ETD dataset. For evaluating the models, we have used both the test split of the synthetic dataset and the full split of the real dataset.
Optimization.
We have used the AdamW optimizer (Loshchilov. & Hutter, 2019) for all tasks and models and trained for 10 epochs. The learning rate has been selected with a random search over the interval , using the validation split of the synthetic ETD dataset. The weight decay values are also randomly searched in the range . The batch size has been tuned over for GRU, for VAP, and for Wav2vec 2.0. The searched hyperparameters are provided in Appendix C.
Hardware.
All experiments have taken place on a single GPU; some experiments have taken place on an NVIDIA L40S, while some have been done on NVIDIA RTX 6000 Ada. To measure the inference latency of ETD models, we used the iPhone 12 mini; the computing took place on the A14 bionic chipset with Hexa-core CPUs.
6 Results
6.1 Binary classification
| Methods | Synthetic data | Real data | ||||||
| Pre. | Rec. | F1 | Acc. | Pre. | Rec. | F1 | Acc. | |
| VAP (Ekstedt & Skantze, 2022) | 48.4 | 48.5 | 48.3 | 57.2 | ||||
| GRU (Chung et al., 2014) | ||||||||
| Wav2vec 2.0 (Baevski et al., 2020) | ||||||||
In Table 4, we report the binary classification results of ETD models. From the table, we observe that the GRU models—which are much lighter than other models—achieve a significantly lower accuracy than other methods, with around 20%p gap in both synthetic and real datasets. This result motivates us to develop a computationally efficient framework that can achieve both efficiency and accuracy.
6.2 Real-time audio segmentation
| Methods | Synthetic data | Real data | ||
| F1 | IoU | F1 | IoU | |
| VAP (Ekstedt & Skantze, 2022) | 92.9 | 87.7 | 17.6 | 10.7 |
| GRU (Chung et al., 2014) | 85.5 | 76.2 | 24.8 | 14.7 |
| Wav2vec 2.0 (Baevski et al., 2020) | 94.7 | 90.2 | 30.3 | 17.9 |
| Ours | 94.0 | 88.9 | 28.0 | 16.4 |
In Table 5, we compare the real-time audio segmentation performances of the proposed SpeculativeETD against three baseline methods. The results demonstrate that the SpeculativeETD achieves a performance that is comparable to a larger-scale model, namely Wav2vec 2.0, less than 2%p difference in IoU. The proposed method also outperforms VAP, which requires much more computation.
6.3 Efficiency of SpeculativeETD
Methods Compute (MFLOPs) VAP 10,354.98 GRU 45.34 Wav2vec 2.0 34,971.68 SpeculativeETD 45.34 + 874.30 = 919.64 Table 6: FLOPs comparison for real-time processing.
FLOPs comparison.
We evaluate the computational cost of our method in terms of FLOPs using 100 synthetic samples, as FLOPs for each sample are detailed in Table 6. By leveraging the server-side Wav2vec 2.0 model only when vital, SpeculativeETD achieves comparable performance to Wav2vec 2.0 with substantially lower FLOPs. Moreover, latency remains unaffected as Wav2vec 2.0 runs on a server GPU rather than on-device.
Methods Latency (ms) Load Init Execute Wav2vec 2.0 874.06 17.89 1500.32 GRU (SpeculativeETD) 1.16 3.85 0.26 Table 7: Latency analysis in a real-time on-device setting. The ‘Execute’ latency is the average inference time when the model performs inference every 100 ms on a 5-second speech audio.
Latency.
We verify whether the latency is suitable for real-time on-device deployment. We measure the latency via LiteRT 111https://ai.google.dev/edge/litert on iPhone mini 12 (iOS 16.3.1). We check model loading, initialization, and execution time, which is shown in Table 7. We averaged three runs of latency measurement, and the execution time was the average of 50 iterations of inference time when the model performed inference every 100 ms on a 5-second speech audio.
Our experiments demonstrate that SpeculativeETD, which utilizes GRU on the on-device side, achieves latency below 1 ms per 100 ms interval. As illustrated in Figure 5, the inference latency of Wav2vec 2.0 significantly increases with input length due to the computational overhead of transformer architectures, while GRU latency remains stable. These results confirm that SpeculativeETD is an excellent option for serving real-time on-device.
7 Conclusion
We introduce the ETD dataset, a novel publicly available dataset for end-turn detection in spoken dialogue systems, and propose SpeculativeETD, an efficient method for real-time end-turn detection. These contributions address critical challenges in developing more natural and responsive spoken dialogue agents, particularly those powered by large language models. The ETD dataset comprises synthetic and real data and is the first public dataset for end-turn detection research. This dataset enables more comprehensive training and evaluation of end-turn detection models by offering a diverse range of conversational scenarios, including artificially extended pauses and injected filler words. Also, real data further enhances the dataset’s utility in assessing model generalizability in real-world conversational settings. SpeculativeETD, represents a significant advancement in balancing efficiency and accuracy for real-time end-turn detection. Combining a lightweight GRU on-device model for on-device processing with a high-performance Wav2Vec 2.0 server-side model on the server improves accuracy while maintaining low latency. This two-stage framework effectively addresses the computational constraints of on-devices while leveraging the superior performance of larger models.
Ethics Statement
All datasets used in this research are exclusively for academic purposes. We release open-source data according to original licensing terms and provide reproducible code for datasets that cannot be directly shared due to licensing constraints.
Appendix A More experiments
Full experiement result
| Methods | Speaking Unit | Gap | Pause | Average | ||||||||||
| Pre. | Rec. | F1 | IoU | Pre. | Rec. | F1 | IoU | Pre. | Rec. | F1 | IoU | F1 Avg. | IoU Avg. | |
| VAP | ||||||||||||||
| GRU | ||||||||||||||
| Wav2vec 2.0 | ||||||||||||||
| Ours | ||||||||||||||
| Methods | Speaking Unit | Gap | Pause | Average | ||||||||||
| Pre. | Rec. | F1 | IoU | Pre. | Rec. | F1 | IoU | Pre. | Rec. | F1 | IoU | F1 Avg. | IoU Avg. | |
| VAP | 51.6 | 35.1 | 41.8 | 26.4 | 2.0 | 6.5 | 3.1 | 1.6 | 10.3 | 6.2 | 7.8 | 4.0 | ||
| GRU | ||||||||||||||
| Wav2vec 2.0 | 55.7 | 23.5 | 33.03 | 19.78 | 21.1 | 54.1 | 30.34 | 17.88 | 28.7 | 26.4 | 27.48 | 15.93 | ||
| Ours | 50.3 | 19.4 | 28.04 | 16.30 | 20.3 | 53.4 | 29.44 | 17.26 | 27.4 | 25.7 | 26.56 | 15.31 | ||
Appendix B Data quality improvement
Segment Filtering.
We applied a structured segmentation process to better capture conversational dynamics. Consecutive speech fragments from the same speaker were merged into a Speaking Unit (SU) if separated by less than 200ms of silence. Silences longer than 200 ms were labeled Pause (if within the same speaker) or Gap (if between different speakers), based on established turn-taking criteria (Stivers et al., 2009; Heldner & Edlund, 2010). Using these labels, we grouped continuous SUs into segments, where each segment represents a stretch of speech bounded by either a Pause or a Gap. To focus on turn-relevant phenomena, we filtered out segments that did not contain any Pause or Gap, retaining only those that reflect actual hesitation or speaker transitions crucial for end-turn detection.
| Language | Code | # Files | Included |
| English | en | 8,022 | ✓ |
| Hindi | hi | 674 | ✗ |
| Romanian | ro | 334 | ✗ |
| Urdu | ur | 315 | ✗ |
| Spanish | es | 258 | ✗ |
| Arabic | ar | 179 | ✗ |
| Swahili | sw | 151 | ✗ |
| Korean | ko | 43 | ✗ |
| Portuguese | pt | 41 | ✗ |
| French | fr | 25 | ✗ |
| Others (40+ languages) | — | 1,055 | ✗ |
Language Filtering.
To ensure the quality and consistency of the real-world speech data, we applied a language filtering step using the Whisper model (Radford et al., 2022). For each audio file, we used Whisper’s language detection capability to identify the spoken language. Since our focus is on English-based dialogue systems, we retained only the samples identified as English and discarded those in other languages. Table 10 summarizes the top 10 most frequent languages and the aggregate of other low-resource languages. Out of a total of 11,127 collected files, 8,022 were classified as English and included in the final dataset. We additionally conducted a human evaluation of language identification performance by Whisper, which is recognized well as English. We checked all the data, and the accuracy of identification was 99.07.
Appendix C Experiment details
| Model | Hyper-parameter | ||
| epochs | batch size | lr | |
| VAP | 10 | 32 | |
| GRU | 10 | 256 | |
| Wav2vec 2.0 | 10 | 8 | |
We detail the hyperparameter settings used for training each model in our experiments. All models were trained for 10 epochs. The VAP model was configured with a batch size of 32 and a learning rate of . For the GRU model, we used a larger batch size of 256 and set the learning rate to . In the case of Wav2vec 2.0, we employed a smaller batch size of 8 with a learning rate of .